[g(x) + (x)](*) [fly)+ (y)] (**) 8(x+h) = g(x)+ fly + k) = f(y)+ (*) gives us f(g(x+h)) =f( (x)+ [8(x) + v(x)])
Answer #1

geath) = gen) + 1 [ge (x) + 4(x+1)] where we are © f(x+x) = 444 +K] [ f(X) + \u7y+k]] where we to using & ③, Then 79 (476))If you have any doubt, please comment and I will try my best to explore this to you.